Find the slope of the line that is a) parallel and b) perpendicular to the given line.

12. x=-8

13. y=10
1
Expert's answer
2013-01-24T07:55:12-0500
x=-8
x+8 = 0
0*y = x+8
y = (1/0)x +8/0
a) k1=1/0 -& undefined slope
b) k1*k2 = 1
k2 = 1/k1 = 0/1 = 0
Answer: a)& undefined slope, b) 0
y = 10
y = 0*x + 10
a) slope - k1 = 0
b) k2 = 1/k1& -& undefined slope
Answer: a) 0,& b)& undefined slope